OpenStudy (anonymous):

3. x = –4y + 4 2x + 8y = 8 (1 point) one two infinite none 4. y = 6x + 2 3y – 18x = 12 (1 point) one two infinite none 5. x – 2y = 6 3x – 6y = 18 (1 point) one two infinite none 6. y – 5x = –6 3y – 15x = –12 (1 point) one two infinite none

OpenStudy (anonymous):

his is solving system of equations right?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

3. x = –4y + 4 2x + 8y = 8 one two infinite none just solve for one variable first x = –4y + 4 this is already solved for x 2x + 8y = 8 next substitute the variable 2 x + 8 y = 8 | V 2 ( –4y + 4) + 8y = 8

OpenStudy (anonymous):

-8y + 8 + 8y = 8 which ends up just 8 = 8

OpenStudy (anonymous):

So its infinite..?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

y = 6x + 2 3y – 18x = 12 this is already solved for y so it is given